The second hand of a clock is four inches long. To the nearest inch, how far does the tip of the second hand travel in 45 seconds?
Answer by nerdybill(7384)

The second hand of a clock is four inches long. To the nearest inch, how far does the tip of the second hand travel in 45 seconds?
.
If the tip went the entire circle of 60 seconds it would have traveled:
Circumference = (pi)(2r)
where
pi is 3.14
r is radius
.
Circumference = (3.14)(2)(4)
Circumference = (3.14)(8)
Circumference = 25.12
.
If we only traveled for 45 seconds:
45/60 * 25.12
= 3/4 * 25.12
= .75 * 25.12
= 18.84
To the nearest inch:
19 inches